# Breaking Boundaries: Investigating the Effects of Model Editing on Cross-linguistic Performance

### v2

@@ Abstract
The integration of pretrained language models (PLMs) like BERT and GPT has revolutionized NLP, particularly for English, but it has also created linguistic imbalances. This paper strategically identifies the need for linguistic equity by examining several knowledge editing techniques in multilingual contexts. We evaluate the performance of models such as Mistral, TowerInstruct, OpenHathi, Tamil-Llama, and Kan-Llama across languages including English, German, French, Italian, Spanish, Hindi, Tamil, and Kannada. Our research identifies significant discrepancies in normal and merged models concerning cross-lingual consistency. We employ strategies like ‘each language for itself’ (ELFI) and ‘each language for others’ (ELFO) to stress-test these models. Our findings demonstrate the potential for LLMs to overcome linguistic barriers, laying the groundwork for future research in achieving linguistic inclusivity in AI technologies  [^1] Repository: https://github.com/NeuralSentinel/BreakingBoundaries.
@@
[ topline=false, bottomline=false, skipabove=skipbelow=leftline=true, rightline=false, linecolor=red, linewidth=4pt, innertopmargin=10pt, innerbottommargin=2pt, innerrightmargin=2pt, innerleftmargin=2pt, backgroundcolor=gray!10, roundcorner=10pt ]stylishframe
@@ 1 Introduction
The advent of model editing techniques, as explored by Sinitsin et al. (2020) and De Cao et al. (2021), introduces an innovative avenue for refining LLM responses to specific inputs, marking a significant stride toward addressing these challenges. However, applying these techniques within a cross-lingual framework introduces a set of unique obstacles Qi et al. (2023); Xu et al. (2023), necessitating the development of inventive strategies to ensure the effective and equitable operation of multilingual PLMs across the globe’s linguistic tapestry Wang et al. (2023).
Building upon this foundation, recent scholarly endeavors have increasingly focused on the nuances of knowledge editing within LLMs, particularly within monolingual contexts Hazra et al. (2024); Banerjee et al. (2024). This body of work has catalyzed significant advancements in the creation and refinement of models tailored for multilingual usage, such as Bloom Workshop et al. (2023), ChatGPT [^2] https://chat.openai.com/, LLaMA Touvron et al. (2023). Yet, the broader implications of implementing these knowledge edits across diverse linguistic landscapes remain an area ripe for exploration.
The exploration of cross-lingual consistency is paramount for several compelling reasons. Primarily, it challenges the conventional wisdom that knowledge is inextricably linked to the form of language, advocating instead for a model’s comprehension to remain steadfast across linguistic divisions. For instance, a model’s recognition of “Dent Island Light, located in: Belgium(Post Edit) (see Figure 1) should be consistent, irrespective of the language employed. Such consistency is crucial for ensuring a uniform user experience across different languages, thereby democratizing access to information and technology.
In pursuit of these objectives, our study conducts a comprehensive examination of the efficacy with which multilingual models transfer knowledge across eight distinct languages tackling the largely unexplored challenges associated with maintaining knowledge consistency and fostering language diversity within the sphere of knowledge editing for LLMs. Inspired by the earlier research Beniwal et al. (2024) in this domain, our investigation explores various strategies for disseminating factual knowledge from linguistically rich to relatively resource-scarce languages, employing the “each language for itself” (ELFI) and “each language for others” (ELFO) principle as a foundational premise Das et al. (2022). Our findings illustrate the intricate dynamics and untapped potential of knowledge transfer in multilingual contexts, underscoring the complexities, challenges, and opportunities that lie ahead in our quest to achieve linguistic equity in the age of artificial intelligence. Our contributions include:
  • We conduct extensive experiments on model editing across eight distinct languages (English (En), German (De), French (Fr), Italian (It) and Spanish (Es from Romance and Germanic scripts, whereas Hindi (Hi), Tamil (Ta) and Kannada (Kn)) from Indic scripts, using two approaches: “each language for itself” (ELFI) and “each language for others” (ELFO). Our observations specifically focused on the performance of decoder-only models from a multilingual perspective.
  • Second, this study explores decoder-only models, including state-of-the-art systems such as Mistral, TowerInstruct, OpenHathi, Tamil-Llama and Kan-Llama (all 7B models). We utilize these models in conjunction with well-known editing methods, ROME and MEMIT, marking a significant advancement in the field.
  • Third, to the best of our knowledge, this paper is the first to evaluate that while model merging enhances capabilities, it still falls short in maintaining cross-lingual consistency post editing.
  • Our comprehensive error analysis reveals significant areas where linguistic discrepancies across different languages lead to varied interpretations and meanings. This study delves into how such differences can cause models to generate incorrect responses post edit.
@@ 2 Related work
Targeted parameter editing specifically focuses on identifying and modifying particular components within a model to integrate new information. This approach was illustrated in Dai et al. (2022) through the identification and adjustment of ‘knowledge neurons’ in transformer models, allowing for the reflection of new data. Meng and colleagues furthered this concept with the Rank-One Model Editing (ROME) Meng et al. (2022) technique, which aimed to update key neural network weights, refreshing factual knowledge in LLMs. Despite their effectiveness for singular updates, these techniques faced difficulties with multiple simultaneous updates. MEMIT Meng et al. (2023), a successor of ROME, addressed this by enabling the concurrent modification of various knowledge points, a breakthrough supported by subsequent studies like Hase et al. (2023); Yao et al. (2023).
Multilingual knowledge editing Research in knowledge editing has predominantly focused on English. However, a handful of studies have ventured into the multilingual domain, translating English-based prompts and object pairs into various languages. Initiatives like X-FACTR Jiang et al. (2020) and M-LAMA Kassner et al. (2021) have identified substantial knowledge disparities across languages, attributed to differences in training data volume. These discrepancies are particularly pronounced for languages other than English and some European languages, where probing accuracies often fall below 10%. GeoMLAMA Yin et al. (2022), took a unique approach by examining region-specific commonsense knowledge, revealing that a country’s native language may not always be the most effective for accessing specific national knowledge.
Our research takes a systematic step in analyzing the cross-lingual consistency of factual knowledge within multilingual LLMs, assessing how consistently an LLM provides the same answer to the same question across different languages. Earlier works, such as those by Wang et al. (2023); Beniwal et al. (2024), initiated this exploration with mBERT, uncovering low consistencies. We expand on this by including various LLMs finetuned on specific languages.
@@ 4 Dataset
For our experiments, we use the popular CounterFact Meng et al. (2022) and ZsRE Levy et al. (2017) datasets. We uniformly sample $\sim$ 550 edit instances from each dataset. Each edit instance in these datasets includes the actual edit case, the reliability prompt, the generalization instances, the locality prompt and its answer, portability and its answer. Further we use google translator  [^3] https://translate.google.com/ to translate each edit instance into seven other languages – German (De), French (Fr), Italian (It), Spanish (Es), Hindi (Hi), Tamil (Ta) and Kannada (Kn). In both the datasets, the actual portability prompt is an interrogative sentence (i.e., in the form of question). However, when the question gets translated to other languages, the translated question becomes different from actual question format. For example, when the actual portability prompt in English “To which language family does the official language of Sastamala belong?” is translated to French the new prompt becomes “À quelle langue la famille appartient la langue officielle de Sastamala?”. However when this is back-translated to English the prompt means “Which family language does the official language of Sastamala belong to?” which is not the same as the original English prompt. We therefore employed GPT-$4$https://openai.com/research/gpt-4, version: gpt-4-0125-preview to convert question in the interrogative sentence into a task of sentence completion. Subsequently we translate this sentence completion form to other languages to obtain the corresponding portability prompt.
@@ 5.1 Selection of LLMs
We use the following multilingual LLMs for our experiments.
@@
Mistral-7B-Instruct-v0.2666https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2: The model was developed by Jiang et al. (2023) and supports multilinguality [^7] https://encord.com/blog/mistral-large-explained/. It is designed around the causal language modeling framework. We shall refer to this model as Mistral.
TowerInstruct-7B-v0.2888https://huggingface.co/Unbabel/TowerInstruct-7B-v0.2: This model Alves et al. (2024) has been developed on top of LLaMA2 Touvron et al. (2023) architecture and supports multilinguality including English, German, French, Spanish, Chinese, Portuguese, Italian, Russian, Korean, and Dutch. We shall refer to this model as TowerInstruct.
@@
Tamil-llama-7b-base-v0.1101010https://huggingface.co/abhinand/tamil-llama-7b-base-v0.1: This is a sophisticated model developed specifically for bilingual tasks in Tamil and English, leveraging a 7 billion parameter causal language modeling framework. We shall refer to this model as Tamil-Llama.
Counterfact Models TowerInstruct Mistral OpenHathi Tamil-Llama Kan-Llama
Languages Metrics ROME MEMIT ROME MEMIT ROME MEMIT ROME MEMIT ROME MEMIT
De Rel 0.83/0.96 0.73/0.83 0.83/0.96 0.73/0.87 - - - - - -
Gen 0.27/0.31 0.19/0.22 0.28/0.31 0.19/0.22 - - - - - -
Loc 0.22/0.23 0.19/0.22 0.21/0.23 0.24/0.27 - - - - - -
Port 0.01/0.01 0.01/0.01 0.03/0.04 0.04/0.06 - - - - - -
Es Rel 0.82/0.92 0.70/0.80 0.81/0.91 0.78/0.86 - - - - - -
Gen 0.33/0.37 0.23/0.27 0.28/0.32 0.22/0.27 - - - - - -
Loc 0.21/0.22 0.19/0.19 0.25/0.27 0.27/0.29 - - - - - -
Port 0.00/0.00 0.00/0.00 0.03/0.03 0.03/0.04 - - - - - -
It Rel 0.87/0.93 0.74/0.78 0.86/0.91 0.80/0.88 - - - - - -
Gen 0.35/0.38 0.25/0.26 0.28/0.30 0.24/0.27 - - - - - -
Loc 0.18/0.19 0.20/0.20 0.26/0.27 0.27/0.28 - - - - - -
Port 0.02/0.02 0.02/0.03 0.02/0.03 0.03/0.03 - - - - - -
Fr Rel 0.83/0.90 0.65/0.72 0.83/0.89 0.79/0.85 - - - - - -
Gen 0.31/0.33 0.22/0.24 0.29/0.30 0.24/0.25 - - - - - -
Loc 0.21/0.22 0.17/0.19 0.20/0.22 0.24/0.25 - - - - - -
Port 0.00/0.01 0.00/0.00 0.03/0.03 0.03/0.03 - - - - - -
Hi Rel - - - - 0.02/0.02 0.45/0.60 - - - -
Gen - - - - 0.00/0.00 0.26/0.33 - - - -
Loc - - - - 0.31/0.35 0.02/0.03 - - - -
Port - - - - 0.01/0.01 0.01/0.01 - - - -
Ta Rel - - - - - - 0.12/0.15 0.48/0.59 - -
Gen - - - - - - 0.03/0.04 0.21/0.25 - -
Loc - - - - - - 0.01/0.01 0.01/0.01 - -
Port - - - - - - 0.01/0.01 0.01/0.01 - -
Kn Rel - - - - - - - - 0.21/0.26 0.14/0.18
Gen - - - - - - - - 0.07/0.08 0.04/0.05
Loc - - - - - - - - 0.03/0.04 0.02/0.03
Port - - - - - - - - 0.00/0.00 0.00/0.01
Table 1: Comparison of reliability, generalization, locality, and portability scores across various language models, evaluated using the CounterFact dataset under same language edit and inference settings. The highest score is presented in bold whereas the second highest is underlined.

### v3

@@ Abstract
Pretrained language models (PLMs) have transformed natural language processing (NLP) but tend to exacerbate linguistic disparities in multilingual contexts. While earlier research has primarily focused on transformer-based models like BERT, this study shifts attention to large language models (LLMs) such as Mistral, TowerInstruct, OpenHathi, Tamil-Llama, and Kan-Llama. Through comprehensive evaluations across eight languages—including high-resource ones (English, German, French, Italian, Spanish) and low-resource ones (Hindi, Tamil, Kannada)—the research uncovers significant shortcomings in ensuring multilingual robustness and adaptability. Employing frameworks like “each language for itself” (ELFI) and “each language for others" (ELFO), the analysis reveals that existing LLMs struggle to address linguistic inequities. Even strategies like model merging fail to close these gaps, highlighting fundamental deficiencies. These findings underscore the urgent need to redesign AI systems to achieve genuine linguistic inclusivity and balanced performance across diverse languages.
@@
[…]
@@ 1 Introduction
Handling multilinguality in language models remains a significant challenge, particularly when models are prompted in languages other than English. Tasks such as question answering Xu et al. (2024), addressing multilingual safety concerns Wang et al. (2024); Deng et al. (2024), or performing knowledge edits Hazra et al. (2024) often reveal noticeable gaps in performance for low-resource languages. Despite the advancements in multilingual large language models (LLMs), disparities persist, especially for languages with fewer computational resources. A clear example of this issue arises in knowledge editing Sinitsin et al. (2020); De Cao et al. (2021). For instance, when an LLM is updated to correct a factual statement, “The PM of the UK is Rishi Sunak" to “The PM of the UK is Keir Starmer" the model may apply the update accurately in well-represented languages like English or French Qi et al. (2023); Xu et al. (2023). However, the same edit often fails to propagate when queried in low-resourced languages like Tamil or Hindi. This inconsistency highlights a critical weakness in the ability of LLMs to transfer factual updates across languages. Even advanced models like Mistral and TowerInstruct, while effective in European languages, struggle significantly with low-resource languages. This limitation undermines the broader goal of making language technologies universally accessible and equitable Wang et al. (2023).
This research aims to uncover the disparities in cross-lingual performance of LLMs to promote future linguistic inclusivity. While model editing techniques have advanced in monolingual settings, ensuring that factual updates made in one language are accurately reflected across others remains a major challenge Hazra et al. (2024); Banerjee et al. (2024). This issue is particularly severe for low-resource languages, where models often fail to maintain reliability and consistency after edits. Such limitations reduce the utility of LLMs for these languages and widen existing linguistic inequities, leaving many communities underserved. Our work highlights these gaps, showing how current models struggle to manage multilingual updates, especially in underrepresented languages. By evaluating cross-lingual performance, we emphasize the need for more inclusive approaches to ensure that LLMs benefit users of all languages, not just those with abundant resources.
In this work, we conduct a comprehensive evaluation of how factual knowledge is transferred and maintained across eight linguistically diverse languages. We examine established knowledge editing techniques such as ROME Meng et al. (2022) and MEMIT Meng et al. (2023) to assess their performance in multilingual contexts. Our research utilizes two strategies Das et al. (2022)—“each language for itself" (ELFI) and “each language for others" (ELFO)—to rigorously test the ability of LLMs to preserve cross-lingual knowledge consistency. Through this evaluation, we reveal current models’ limitations in maintaining consistent cross-lingual edits, emphasizing critical gaps to address for enhancing LLMs, particularly in low-resource languages. Our key contributions are as follows.
[…]
@@ 2 Related work
Targeted parameter editing modifies specific model components to integrate new information. Dai et al. (2022) introduced adjustments to ‘knowledge neurons’ in transformers, while ROME Meng et al. (2022) updated neural weights to refresh LLM knowledge. MEMIT Meng et al. (2023) expanded ROME for simultaneous updates, with further validation by Hase et al. (2023); Yao et al. (2023).
Multilingual knowledge editing remains limited, focusing mainly on translating English prompts. X-FACTR Jiang et al. (2020) and M-LAMA Kassner et al. (2021) exposed large knowledge gaps in non-English languages, often with $<10\%$ accuracy. GeoMLAMA Yin et al. (2022) revealed that native languages may not best access national knowledge. We analyze cross-lingual consistency in multilingual LLMs, extending prior work mostly on BERT (pre LLM era) to diverse LLMs fine-tuned for specific languages Wang et al. (2023); Beniwal et al. (2024).
@@ 4 Dataset
For our experiments, we use the popular CounterFact Meng et al. (2022) and ZsRE Levy et al. (2017) datasets. We uniformly sample $\sim$ 550 edit instances from each dataset. Each edit instance in these datasets includes the actual edit case, the reliability prompt, the generalization instances, the locality prompt and its answer, portability and its answer. Further we use google translator  [^1] https://translate.google.com/ to translate each edit instance into seven other languages – German (De), French (Fr), Italian (It), Spanish (Es), Hindi (Hi), Tamil (Ta) and Kannada (Kn). In both the datasets, the actual portability prompt is an interrogative sentence (i.e., in the form of question). However, when the question gets translated to other languages, the translated question becomes different from actual question format. For example, when the actual portability prompt in English “To which language family does the official language of Sastamala belong?” is translated to French the new prompt becomes “À quelle langue la famille appartient la langue officielle de Sastamala?”. However when this is back-translated to English the prompt means “Which family language does the official language of Sastamala belong to?” which is not the same as the original English prompt. We therefore employed GPT-4222openai.com/research/gpt-4, version: gpt-4-0125-preview to convert question in the interrogative sentence into a task of sentence completion. Subsequently we translate this sentence completion form to other languages to obtain the corresponding portability prompt.
@@ 5.1 Selection of LLMs
We use the following multilingual LLMs for our experiments:
@@
Mistral-7B-Instruct-v0.2 (Mistral)444huggingface.co/mistralai/Mistral-7B-Instruct-v0.2: A multilingual causal language model Jiang et al. (2023), supporting diverse languages [^5] https://encord.com/blog/mistral-large-explained/.
TowerInstruct-7B-v0.2 (TowerInstruct)666huggingface.co/Unbabel/TowerInstruct-7B-v0.2: Based on LLaMA2 Touvron et al. (2023), supports multilinguality across 10 languages, including English, German, and Chinese.
OpenHathi-7B-Hi-v0.1-Base (OpenHathi)777huggingface.co/sarvamai/OpenHathi-7B-Hi-v0.1-Base: Optimized for Indian languages like Hindi and Tamil using a GPT-3-like transformer with hybrid partitioned attention.
@@
Tamil-llama-7b-base-v0.1 (Tamil-Llama)888huggingface.co/abhinand/tamil-llama-7b-base-v0.1: A bilingual Tamil-English model Balachandran (2023) using a 7B-parameter causal language framework.
Kan-LLaMA-7B-SFT (Kan-Llama)999huggingface.co/Tensoic/Kan-Llama-7B-SFT-v0.5: Specialized in Kannada with a 49,420-token vocabulary, pre-trained on 600M tokens from CulturaX using low-rank adaptation. More details on models are in Appendix A.
Languages CounterFact ZsRE
Models TowerInstruct Mistral TowerInstruct Mistral
Metrics RO ME RO ME RO ME RO ME
De Rel 0.83/0.96 0.73/0.83 0.83/0.96 0.73/0.87 0.48/0.59 0.25/0.30 0.51/0.62 0.38/0.47
Gen 0.27/0.31 0.19/0.22 0.28/0.31 0.19/0.22 0.33/0.39 0.11/0.12 0.35/0.45 0.18/0.24
Loc 0.22/0.23 0.19/0.22 0.21/0.23 0.24/0.27 0.00/0.01 0.00/0.01 0.01/0.02 0.01/0.03
Port 0.01/0.01 0.01/0.01 0.03/0.04 0.04/0.06 0.02/0.02 0.00/0.00 0.08/0.10 0.02/0.04
Es Rel 0.82/0.92 0.70/0.80 0.81/0.91 0.78/0.86 0.44/0.59 0.24/0.34 0.49/0.61 0.37/0.49
Gen 0.33/0.37 0.23/0.27 0.28/0.32 0.22/0.27 0.30/0.40 0.16/0.20 0.35/0.45 0.22/0.29
Loc 0.21/0.22 0.19/0.19 0.25/0.27 0.27/0.29 0.00/0.01 0.01/0.02 0.01/0.01 0.02/0.02
Port 0.00/0.00 0.00/0.00 0.03/0.03 0.03/0.04 0.02/0.02 0.01/0.02 0.03/0.07 0.03/0.04
It Rel 0.87/0.93 0.74/0.78 0.86/0.91 0.80/0.88 0.54/0.62 0.25/0.29 0.58/0.65 0.42/0.50
Gen 0.35/0.38 0.25/0.26 0.28/0.30 0.24/0.27 0.35/0.43 0.16/0.20 0.42/0.48 0.25/0.31
Loc 0.18/0.19 0.20/0.20 0.26/0.27 0.27/0.28 0.00/0.00 0.00/0.01 0.00/0.02 0.01/0.02
Port 0.02/0.02 0.02/0.03 0.02/0.03 0.03/0.03 0.01/0.02 0.02/0.03 0.07/0.08 0.01/0.03
Fr Rel 0.83/0.90 0.65/0.72 0.83/0.89 0.79/0.85 0.51/0.59 0.27/0.35 0.52/0.63 0.40/0.50
Gen 0.31/0.33 0.22/0.24 0.29/0.30 0.24/0.25 0.28/0.35 0.14/0.17 0.40/0.50 0.19/0.27
Loc 0.21/0.22 0.17/0.19 0.20/0.22 0.24/0.25 0.00/0.01 0.00/0.02 0.01/0.02 0.01/0.02
Port 0.00/0.01 0.00/0.00 0.03/0.03 0.03/0.03 0.03/0.05 0.03/0.03 0.06/0.09 0.04/0.06
Table 1: Comparison of reliability, generalization, locality, and portability scores across language models under Self edit - self inference settings. The highest scores for individual metrics in ROME and MEMIT are highlighted in magenta for CounterFact and in cyan for ZSRE, with values shown as Exact Match/Partial Match.
Languages CounterFact ZsRE
Models TowerInstruct Mistral TowerInstruct Mistral
Metrics RO ME RO ME RO ME RO ME
De Rel 0.48/0.53 0.40/0.46 0.50/0.56 0.54/0.61 0.24/0.28 0.10/0.14 0.34/0.45 0.14/0.18
Gen 0.25/0.27 0.13/0.17 0.23/0.27 0.22/0.23 0.18/0.23 0.12/0.14 0.26/0.35 0.14/0.16
Loc 0.20/0.21 0.19/0.22 0.23/0.25 0.26/0.28 0.00/0.01 0.00/0.02 0.01/0.02 0.01/0.03
Port 0.00/0.00 0.00/0.00 0.03/0.03 0.03/0.04 0.02/0.02 0.02/0.02 0.06/0.07 0.02/0.03
Es Rel 0.51/0.56 0.40/0.48 0.57/0.62 0.56/0.60 0.24/0.29 0.12/0.14 0.39/0.48 0.19/0.26
Gen 0.26/0.29 0.18/0.22 0.25/0.29 0.21/0.26 0.18/0.25 0.09/0.11 0.33/0.41 0.14/0.21
Loc 0.22/0.24 0.17/0.17 0.24/0.27 0.25/0.27 0.00/0.01 0.01/0.02 0.01/0.02 0.02/0.02
Port 0.00/0.00 0.00/0.00 0.03/0.03 0.03/0.04 0.02/0.03 0.01/0.01 0.04/0.06 0.04/0.05
It Rel 0.45/0.50 0.35/0.40 0.47/0.58 0.44/0.49 0.24/0.29 0.12/0.14 0.31/0.34 0.23/0.27
Gen 0.23/0.27 0.19/0.20 0.25/0.35 0.21/0.23 0.17/0.22 0.11/0.13 0.26/0.32 0.18/0.21
Loc 0.20/0.21 0.20/0.20 0.24/0.36 0.28/0.29 0.00/0.00 0.00/0.01 0.00/0.02 0.01/0.02
Port 0.01/0.02 0.01/0.02 0.03/0.11 0.04/0.04 0.01/0.02 0.02/0.02 0.07/0.08 0.01/0.01
Fr Rel 0.50/0.53 0.45/0.49 0.49/0.55 0.51/0.59 0.22/0.26 0.12/0.17 0.36/0.44 0.23/0.28
Gen 0.28/0.31 0.19/0.22 0.28/0.31 0.26/0.27 0.15/0.21 0.08/0.10 0.29/0.33 0.16/0.21
Loc 0.23/0.23 0.19/0.21 0.20/0.36 0.25/0.26 0.00/0.01 0.00/0.02 0.01/0.03 0.01/0.02
Port 0.01/0.01 0.01/0.01 0.01/0.12 0.03/0.04 0.02/0.02 0.02/0.02 0.06/0.09 0.04/0.05
Table 2: Comparison of reliability, generalization, locality, and portability scores across language models under English edit - self inference settings. The highest scores for individual metrics in ROME and MEMIT are highlighted in magenta for CounterFact and in cyan for ZSRE, with values shown as Exact Match/Partial Match.
Languages/ Models self edit - self inference (English edit - self inference)
CounterFact ZsRE CounterFact ZsRE
Metrics RO ME RO ME RO ME RO ME
Hi/ OpenHathi Rel 0.02/0.02 0.45/0.60 0.03/0.06 0.20/0.33 0.56/0.66 0.02/0.03 0.03/0.03 0.03/0.06
Gen 0.00/0.00 0.26/0.33 0.01/0.04 0.19/0.28 0.27/0.34 0.03/0.03 0.03/0.03 0.04/0.08
Loc 0.31/0.35 0.02/0.03 0.01/0.01 0.00/0.01 0.26/0.31 0.03/0.03 0.00/0.00 0.00/0.01
Port 0.01/0.01 0.01/0.01 0.00/0.00 0.03/0.03 0.02/0.02 0.00/0.01 0.00/0.00 0.01/0.01
Ta/ Tamil-Llama Rel 0.12/0.15 0.48/0.59 0.06/0.08 0.16/0.21 0.00/0.00 0.01/0.01 0.00/0.00 0.01/0.01
Gen 0.03/0.04 0.21/0.25 0.03/0.04 0.10/0.14 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00
Loc 0.01/0.01 0.01/0.01 0.00/0.00 0.00/0.00 0.01/0.01 0.01/0.02 0.00/0.00 0.00/0.00
Port 0.01/0.01 0.01/0.01 0.00/0.00 0.01/0.01 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00
Kn/ Kan-Llama Rel 0.21/0.26 0.14/0.18 0.16/0.21 0.05/0.07 0.01/0.01 0.00/0.00 0.00/0.01 0.00/0.01
Gen 0.07/0.08 0.04/0.05 0.08/0.17 0.05/0.05 0.00/0.01 0.00/0.00 0.00/0.00 0.00/0.00
Loc 0.03/0.04 0.02/0.03 0.00/0.00 0.00/0.00 0.02/0.02 0.03/0.03 0.00/0.00 0.00/0.00
Port 0.00/0.00 0.00/0.01 0.00/0.01 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00 0.00/0.00
Table 3: Comparison of scores in indic language models. Highest scores are in bold, second-highest underlined, with values shown as Exact Match/Partial Match.